# Analyzing Multifiltering Functions Using Multiparameter Discrete Morse Theory

### v1

@@ 3. $\mathrm{(3)}$
That being said, previous algorithms that generate discrete gradient fields outperform GenerateMDM in terms of speed. In particular, ComputeDiscreteGradient computes a gradient field compatible with the max-extension of a component-wise injective vertex map in $O(|K|+|K_{0}|\log|K_{0}|)$ time, assuming $\dim K$ is low and the star of each vertex is small. However, under the same hypotheses and assuming $k$ is also small, GenerateMDM takes $O(|K|+|f(K)|\log|f(K)|)$ time where $|K_{0}|\leq|f(K)|\leq|K|$. Hence, at best, it can be as efficient as ComputeDiscreteGradient, but it is definitely slower when the number $|f(K)|$ of level sets of $f$ is great. Nonetheless, as mentioned above, function GenerateMDM proves itself to be advantageous on many other levels.
5. Proof of correctness
We now verify that our algorithms produce the desired result. To do so, we first show that every $\sigma\in K$ is processed exactly once. From Algorithm 1, it is easy to see that each level set $L_{u}$ is used exactly once as a parameter in ExpandMDM. Thus, it suffices to show that, for a given $L_{u}$, each $\sigma\in L_{u}$ is processed exactly once by ExpandMDM.
Lemma 5.1.
Proof.
First, suppose $\sigma$ was processed as critical at line 26 of Algorithm 2. Then, $\sigma$ was in PQzero at some point, so all its facets have been processed and $\texttt{processed}(\alpha)=\textbf{True}$. In the case where $\sigma$ is paired with a facet $\tau<\sigma$, then $\sigma$ was popped from PQone at line 12 and $\tau=\texttt{unprocessed\_facet}(\sigma,L_{u})$. For all $\alpha\neq\tau$, we have $\texttt{processed}(\alpha)=\textbf{True}$ because $\tau$ is the unique unprocessed facet of $\sigma$. For $\alpha=\tau$, then $\alpha$ is paired with $\sigma$ which are both declared as processed simultaneously at line 19.
[…]
Now, assume $\sigma$ is the $n^{\text{th}}$ simplex in $L_{u}$ to be paired with a cofacet $\sigma^{\prime}$ and suppose that when the first $n-1$ such simplices had no unprocessed facet in $L_{u}$ when they were paired. Consider a facet $\alpha<\sigma$ in $L_{u}$. As above, we could show that there exists a processed facet $\gamma\in L_{u}$ of $\sigma$ such that $\gamma\neq\sigma$ and $\alpha<\gamma<\sigma^{\prime}$.
Lemma 5.2.
Proof.
[…]
From these observations, we can summarize the process of a single simplex $\sigma\in L_{u}$ with the flowchart in Figure 6. Although it contains a few loops, we can show $\sigma$ always ends up being processed.
Proposition 5.3.
For all $u\in f(K)$, each simplex in $L_{u}$ is processed exactly once by Algorithm 2.
Proof.
The result follows from the three following statements:
  1. $\mathrm{(1)}$
    Each simplex in $L_{u}$ eventually enters PQone or PQzero.
  2. $\mathrm{(2)}$
    Every simplex in PQone and PQzero is eventually processed.
  3. $\mathrm{(3)}$
    A simplex that has already been processed cannot enter PQone or PQzero again.
  • At lines 7 and 9, all simplices of $L_{u}$ are still unprocessed.
  • At lines 20 and 27, a simplex $\sigma$ can only be added to PQone by add_cofacets after one of its facet is processed. However, when $\texttt{processed}(\sigma)=\textbf{True}$, we know from Lemma 5.1 that all facets of $\sigma$ are processed, so add_cofacets cannot add $\sigma$ to PQone.
    • if $\sigma$ does not have priority in PQone, another simplex is popped from PQone is either moved to PQzero or paired with a facet $\tau\in\texttt{PQzero}$ from Lemma 5.2, so $\sigma$ remains unprocessed.
  • $L_{u}\subset K$ is finite;
  • simplices in PQzero cannot enter PQone;
  • processed simplices cannot return to either PQone or PQzero;
[…]
Thus, since Proposition 5.3 stands for every level set $L_{u}\subset K$ and each $L_{u}$ is processed exactly once in Algorithm 1, we deduce that every $\sigma\in K$ is processed exactly once by Algorithm 1. We now show that the dictionaries $g$ and ${\mathcal{V}}$ produced by GenerateMDM indeed represent a mdm function $g:K\rightarrow{\mathbb{R}}^{k}$ and its gradient field ${\mathcal{V}}:K\nrightarrow K$. To do so, a few preliminary observations are needed.
Lemma 5.4.
Proof.
Consider $u\in f(K)$ such that $\sigma\in L_{u}$. For all $\tau<\sigma$, we know that $f(\tau)\preceq f(\sigma)=u$, so either $f(\tau)=u^{\prime}\precneqq u$ or $f(\tau)=u$.
  • If $f(\tau)=u$, $\tau$ was processed within the same level set $L_{u}$ as $\sigma$, and we know from Lemma 5.1 that $\tau$ cannot be processed before $\sigma$. ∎
Lemma 5.5.
Proof.
Consider $u\in f(K)$ such that $\sigma_{0}\in L_{u}$. Since $\sigma_{0}$ is the first simplex to be processed, it belongs to the first level set $L_{u}$ given by the LevelSets routine, which implies that there exists no level set $L_{u^{\prime}}$ with $u^{\prime}\precneqq u$. Now, suppose $\sigma_{0}$ is processed as either one of the simplices $\sigma$ or $\tau$ from lines 16-20. From the properties of a simplicial complex, we know $\sigma$ has at least another facet $\alpha\neq\tau$ and, since $f(\alpha)\preceq f(\sigma)=u$ and there exists no $u^{\prime}\precneqq u$, it follows that $\alpha\in L_{u}$. However, $\tau$ is the unique unprocessed facet of $\sigma$ in $L_{u}$, so $\texttt{processed}(\alpha)=\textbf{True}$, which is a contradiction since $\sigma_{0}\neq\alpha$ and $\sigma_{0}$ is the first simplex to be processed.
Proposition 5.6.
Consider function ComputeG described in Algorithm 3.
  1. $\mathrm{(1)}$
    When called in Algorithm 2, ComputeG always returns a value $u\in{\mathbb{R}}^{k}$ without error, so the output dictionary $g$ of Algorithm 1 is such that $g(\sigma)$ is well defined for all $\sigma\in K$.
  2. $\mathrm{(2)}$
  3. $\mathrm{(3)}$
    For all $\sigma\in K$, we have
    $g_{1}(\sigma)$ $=f_{1}(\sigma)+m\delta\text{ for some }m\in{\mathbb{N}},$
    $g_{i}(\sigma)$ $=f_{i}(\sigma)\text{ for each }i=2,...,k$
Proof.
For $n=0$, consider $u\in f(K)$ such that $\sigma_{0}\in L_{u}$. We know from Lemma 5.5 that $\sigma_{0}$ is a critical vertex, so $g(\sigma_{0})$ is defined by calling $\texttt{ComputeG}(f,g,\delta,\sigma_{0})$ at line 24 of Algorithm 2. From Algorithm 3, it is obvious that $\texttt{ComputeG}(f,g,\delta,\sigma_{0})$ returns $f(\sigma_{0})\in{\mathbb{R}}^{k}$, so $g(\sigma_{0}):=f(\sigma_{0})$, which proves both $\mathrm{(1)}$ and $\mathrm{(3)}$. Also, because $\sigma_{0}$ is a vertex, it has no facet, which makes $\mathrm{(2)}$ trivial.
Now, consider $n\geq 1$ and suppose the proposition is verified for $\sigma_{0},\sigma_{1},...,\sigma_{n-1}$. If $\sigma_{n}$ is a critical vertex, we prove the proposition as above in the case $n=0$. Also, when $\sigma_{n}$ is processed as critical but not a vertex, the proof is similar to the case when $\sigma_{n}$ is paired with a facet by ExpandMDM. Thus, for the remainder of the proof, we assume $\sigma_{n}$ is processed as part of a pair $\tau<\sigma$. In this case, note that $g(\sigma_{n})$ is defined by $\texttt{ComputeG}(f,g,\delta,\sigma,\tau)$ at line 17 of Algorithm 2 and $\sigma$ is not a vertex, so ComputeG necessarily enters the else statement at line 5.
  • $\mathrm{(1)}$
    Consider the set $A$ of facets $\alpha<\sigma$ such that $\alpha\neq\tau$. Every $\alpha\in A$ was processed before $\sigma$ by Lemma 5.4, so $g(\alpha)$ is well defined for each $\alpha\in A$ by the induction hypothesis. Hence, the instruction at line 7 and the conditional statement at line 10 of Algorithm 3 are executed without error, which proves $\mathrm{(1)}$.
  • $\mathrm{(2)}$
    • We first assume $\sigma_{n}=\sigma$. Consider the set $A$ of facets $\alpha<\sigma$ such that $\alpha\neq\tau$ and the vector value $u=(u_{1},...,u_{k})\in{\mathbb{R}}^{k}$ as defined at line 9 of Algorithm 3, so $u_{1}=\max\left(\{f_{1}(\sigma)\}\cup\{g_{1}(\alpha)\ |\ \alpha\in A\}\right)$ and $u_{i}=f_{i}(\sigma)$ for $i>1$. For $i=1$, it is clear that $u_{1}\geq g_{1}(\alpha)$ for all $\alpha\in A$, while for $i>1$,
      $u_{i}=f_{i}(\sigma)\geq f_{i}(\alpha)=g_{i}(\alpha)\ \forall\alpha\in A$
      because $f(\sigma)\succeq f(\alpha)$ by definition of an admissible map and $f_{i}(\alpha)=g_{i}(\alpha)$ for all $i>1$ and all $\alpha\in A$ by the induction hypothesis. Thus, $u\succeq g(\alpha)$ for all $\alpha\in A$. Since $g_{1}(\sigma)$ is defined either as $u_{1}$ or $u_{1}+\delta$ and $g_{i}(\sigma):=u_{i}$ for $i>1$, we have $g(\sigma)\succeq u\succeq g(\alpha)$ for all $\alpha\in A$. Furthermore, if $u\succneqq g(\alpha)$ for all $\alpha\in A$, then $g(\sigma):=u\succneqq g(\alpha)$ for all $\alpha\in A$, and if $u=g(\alpha)$ for some $\alpha\in A$, then $g_{1}(\sigma):=u_{1}+\delta>u_{1}\geq g_{1}(\alpha)$. In both cases, we see that $g(\sigma)\succneqq g(\alpha)$ for all $\alpha\in A$, so $g(\sigma)=g(\alpha)$ for some facet $\alpha<\sigma$ implies $\alpha\notin A$, and thus $\alpha=\tau$.
      Recall from line 17 of Algorithm 2 that $g(\tau)=g(\sigma)$. Moreover, we know from the properties of a simplicial complex that there exists a simplex $\alpha\neq\tau$ such that $\gamma<\alpha<\sigma$. Using the arguments from the previous case, we can see that $g(\sigma)\succneqq g(\alpha)$ since $\alpha$ is a facet of $\sigma$ such that $\alpha\neq\tau$. Also, $\alpha$ was processed before $\sigma$ and $\tau$ from Lemma 5.4, so we can deduce from the induction hypothesis that the facet $\gamma$ of $\alpha$ is such that $g(\alpha)\succeq g(\gamma)$. We conclude that
      $g(\tau)=g(\sigma)\succneqq g(\alpha)\succeq g(\gamma).$
  • $\mathrm{(3)}$
    Recall that $g_{1}(\sigma_{n})$ is either $u_{1}$ or $u_{1}+\delta$ where $u_{1}$ is defined at line 7 of Algorithm 3 as $u_{1}=\max\left(\{f_{1}(\sigma)\}\cup\{g_{1}(\alpha)\ |\ \alpha\in A\}\right)$ and $A$ is the set of facets $\alpha<\sigma$ such that $\alpha\neq\tau$. Also, we see that $f_{1}(\sigma)=f_{1}(\tau)=f_{1}(\sigma_{n})$ since $\sigma_{n}$ is either $\sigma$ or $\tau$ and they both belong in a same level set.
    • Assume $u_{1}=f_{1}(\sigma)$. Then, either $g_{1}(\sigma_{n}):=f_{1}(\sigma_{n})$ or $g_{1}(\sigma_{n}):=f_{1}(\sigma_{n})+\delta$, so $g_{1}(\sigma_{n})=f_{1}(\sigma_{n})+m\delta$ where $m\in\{0,1\}$. Since $n\geq 1$ by hypothesis, we have $m\leq n$.
    • Assume $u_{1}=g_{1}(\alpha)>f_{1}(\sigma)$ for some $\alpha\in A$. By the induction hypothesis, $g_{1}(\alpha)=f_{1}(\alpha)+m\delta$ where $m$ is bounded by the number of simplices processed before $\alpha$, so $u_{1}=f_{1}(\alpha)+m\delta$. Necessarily, $m\leq n-1$. Also, we know that $f_{1}(\alpha)\leq f_{1}(\sigma)$ because $f$ is an admissible map.
        If $f_{1}(\alpha)=f_{1}(\sigma)$, since $f_{1}(\sigma)=f_{1}(\sigma_{n})$ and $g_{1}(\sigma_{n})$ equals either $u_{1}$ or $u_{1}+\delta$, we then have $g_{1}(\sigma_{n})=f_{1}(\sigma_{n})+m\delta$ or $g_{1}(\sigma_{n})=f_{1}(\sigma_{n})+(m+1)\delta$ where both $m,m+1\leq n$, hence the result.
        We show the inequality $f_{1}(\alpha)<f_{1}(\sigma)$ leads to a contradiction. Recall from line 7 of Algorithm 1 that $0<\delta\leq\frac{f_{1}(\sigma)-f_{1}(\alpha)}{|K|}$. Also, $m<n<|K|$, so $m\delta<|K|\delta\leq f_{1}(\sigma)-f_{1}(\alpha)$. Therefore, $g_{1}(\alpha)=f_{1}(\alpha)+m\delta<f_{1}(\sigma)$, which contradicts the assumption that $u_{1}=g_{1}(\alpha)>f_{1}(\sigma)$. ∎
We are now ready to prove GenerateMDM produces the desired result.
Theorem 5.7.
For any simplicial complex $K$, multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$, admissible indexing map $I:K\rightarrow{\mathbb{R}}$ and $\epsilon>0$ given as input GenerateMDM, the outputs $g$ and ${\mathcal{V}}$ correspond to a $f$-compatible mdm function $g:K\rightarrow{\mathbb{R}}^{k}$ and its gradient field ${\mathcal{V}}:K\nrightarrow K$ such that, for all $\sigma\in K$,
$\lVert g(\sigma)-f(\sigma)\rVert<\epsilon$
[…]
Proof.
since $\delta$ is defined in Algorithm 1 so that $\delta\leq\frac{\epsilon}{|K|}$. Next, we show $g$ is a mdm function and ${\mathcal{V}}$ is its gradient field. Let $\sigma\in K$.
We have finally proved that the proposed algorithms produce the desired output. In the next proposition, we highlight another interesting result about the mdm function produced by GenerateMDM.
Proposition 5.8.
Let $g:K\rightarrow{\mathbb{R}}^{k}$ be the mdm function produced by GenerateMDM when given $f:K\rightarrow{\mathbb{R}}^{k}$ as input. The function $f$ is itself mdm if and only if $f=g$.
Proof.
If $f=g$, since $g$ is mdm by Theorem 5.7, we obviously have $f$ mdm as well. Thus, we suppose $f$ is mdm and show that $f(\sigma)=g(\sigma)$ for all $\sigma\in K$.
As in the proof of Proposition 5.6, we label the simplices $\{\sigma_{0},\sigma_{1},...,\sigma_{n},...,\sigma_{N}\}$ of $K$ so that $n<n^{\prime}$ when $\sigma_{n}$ is processed before $\sigma_{n^{\prime}}$ by GenerateMDM and prove all our claim by induction on $n$.
Since $\sigma_{0}$ is a vertex processed as critical by Lemma 5.5, we have that $g(\sigma_{0})$ is defined as $f(\sigma_{0})$, as seen at line 4 of Algorithm 3. For $n\geq 1$, if $\sigma_{n}$ is a vertex and processed as critical, we use the same argument to prove the result. Otherwise, $g(\sigma_{n})$ is defined by going through lines 511 of Algorithm 3 either by processing $\sigma_{n}$ as a critical simplex $\sigma$ or as part of a pair $\tau<\sigma$. In both cases, we know from the induction hypothesis that $g_{1}(\alpha)=f_{1}(\alpha)$ for all $\alpha\in A=\{\text{facets }\alpha<\sigma\ |\ \alpha\neq\tau\}$. Hence, $u_{1}$ is defined at line 7 as
$u_{1}:=\max\left(\{f_{1}(\sigma)\}\bigcup_{\alpha\in A}\{g_{1}(\alpha)\}\right)=\max\left(\{f_{1}(\sigma)\}\bigcup_{\alpha\in A}\{f_{1}(\alpha)\}\right)=f_{1}(\sigma)$
since $f(\sigma)\succeq f(\alpha)$ by definition of an admissible map $f$. Therefore, at line 9, $u$ is defined as $f(\sigma)$. If ComputeG does not enter the if statement at line 10, then $u=f(\sigma)$ is returned to define $g(\sigma)$ and we have the result. Thus, all that is left to prove is that the conditional statement is never verified, i.e. we show $u\neq g(\alpha)$, or equivalently $f(\sigma)\neq f(\alpha)$, for all $\alpha\in A$.
Since $f$ is mdm, we know we have either $\operatorname{card}\mathcal{T}_{f}(\sigma)=0$ or $\operatorname{card}\mathcal{T}_{f}(\sigma)=1$. When $\operatorname{card}\mathcal{T}_{f}(\sigma)=0$, we have directly $f(\sigma)\succneqq f(\alpha)$ for all facets $\alpha<\sigma$ by definition of a mdm function. When $\operatorname{card}\mathcal{T}_{f}(\sigma)=1$, we could show that the $\sigma$ is paired by ExpandMDM with its unique facet in $\mathcal{T}_{f}(\sigma)$, so $\mathcal{T}_{f}(\sigma)=\{\tau\}$. Again, we deduce that for all facets $\alpha\in A$, namely all facets $\alpha\neq\tau$, we have $f(\sigma)\succneqq f(\alpha)$. ∎
@@ 6.1. Defining critical simplices for a multifiltering function
Consider a combinatorial gradient field ${\mathcal{V}}$ compatible with a multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$. As stated in Section 3.2, we know that for any $u\in f(K)$, the number $m_{p}(u)$ of critical $p$-simplices of ${\mathcal{V}}$ in $L_{u}$ is at least $\operatorname{rank}H_{p}(K(u),\bigcup_{u^{\prime}\precneqq u}K(u^{\prime}))$. This expression can be rewritten in terms of $L_{u}$. Indeed, notice that $\bigcup_{u^{\prime}\precneqq u}K(u^{\prime})$ is simply $K(u)\backslash L_{u}$. Then, using the excision theorem (see [26]) to excise $K(u)\backslash\operatorname{Cl}L_{u}$, we see that
@@
Let ${\mathcal{V}}$ be combinatorial vector field compatible with a multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$. For all $u\in f(K)$, if $H_{*}\big{(}\operatorname{Cl}L_{u},\operatorname{Ex}L_{u}\big{)}$ is nonzero, then ${\mathcal{V}}$ has a critical simplex in $L_{u}$. Conversely, if ${\mathcal{V}}$ is perfect relatively to $f$ and has a critical simplex in $L_{u}$, then $H_{*}\big{(}\operatorname{Cl}L_{u},\operatorname{Ex}L_{u}\big{)}$ is nonzero.
@@
Definition 6.2 (Pareto set of multifiltering function).
For a multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$, we say $u\in f(K)$ is a Pareto critical value of $f$ if the relative homology $H_{*}(\operatorname{Cl}L_{u},\operatorname{Ex}L_{u})$ is nonzero. Similarly, $\sigma\in K$ is Pareto critical for $f$ if it belongs in a connected component $C$ of $L_{f(\sigma)}$ such that $H_{*}(\operatorname{Cl}C,\operatorname{Ex}C)$ is nonzero. The Pareto critical simplices of $f$ constitute the Pareto set of $f$, noted $\mathcal{P}_{f}$.
Remark.
@@
Our definition is also inspired from the theory of combinatorial multivector fields, which provides discrete analogues to many concepts from smooth dynamics [51]. Indeed, consider a multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$ and some $u\in f(K)$. For any two simplices $\sigma\leq\sigma^{\prime}$, if $f(\sigma)=f(\sigma^{\prime})=u$, then for all $\tau\in K$ such that $\sigma\leq\tau\leq\sigma^{\prime}$, we have $f(\tau)=u$ because $f(\sigma)\preceq f(\tau)\preceq f(\sigma^{\prime})$ by the definition of a multifiltering function. Hence, $\sigma,\sigma^{\prime}\in L_{u}$ implies $\tau\in L_{u}$ whenever $\sigma\leq\tau\leq\sigma^{\prime}$, meaning that $L_{u}$ is convex with respect to the face relation in $K$. In other words, each $L_{u}$ is a combinatorial multivector in the sense of [51]. Therefore, the partition of $K$ in level sets is a well-defined combinatorial multivector field for which each $L_{u}$ is a multivector. Finally, a multivector $L_{u}$ is defined as critical by Lipiński et al. [51] when $H_{*}(\operatorname{Cl}L_{u},\operatorname{Ex}L_{u})$ is nonzero, so our proposed definition of a Pareto critical value agrees with the combinatorial multivector field theory.
@@
Figure 7. In each figure is shown a neighbourhood $U$ of some point $p$ on the sphere $S^{2}$. For $f:S^{2}\rightarrow{\mathbb{R}}^{2}$ which maps each point to its coordinates $(x,y)\in{\mathbb{R}}^{2}$, the Pareto set of $f$ consists of the two red arcs, while points $q\in U$ such that $f(q)\preceq f(p)$ are orange, those for which $f(p)\preceq f(q)$ are green and all others are yellow. In 7a, the considered point $p$ is regular. In 7b, $p$ is a Pareto minimum and in 7c, it is a Pareto maximum.
@@
To further justify the proposed definition of Pareto singularity, we present concisely the idea of Pareto points for smooth maps and give an interpretation which is common to both discrete and smooth settings. The reader is referred to [46] for more details regarding the basic concepts of differential geometry. Let $f=(f_{1},...,f_{k}):M\rightarrow{\mathbb{R}}^{k}$ be a smooth map on a manifold $M$ and consider some point $p\in M$. We say $p$ is in the Pareto set $\theta$ of $f$ if there exists no direction in which the gradient vectors of all components $f_{i}$ at $p$ agree. In other words, away from $\theta$, all $f_{i}$ increase in some common direction. For all $p\notin\theta$ and for some small enough neighbourhood $U$ of $p$, we could show that this common direction may be represented by the subset $\{q\in U\ |\ f_{i}(q)>f_{i}(p)\text{ for each }i=1,...,k\}$, which is homeomorphic to the intersection of a convex cone with some small neighbourhood, so it is a contractible subset of $U$. An example of such a subset is represented in green in Figure 7a. Generally, this is not true for points $p\in\theta$. For example, in Figure 7b, we see that the subset $\{q\in U\ |\ f_{i}(q)>f_{i}(p)\text{ for each }i=1,...,k\}$ has two distinct components while, in Figure 7c, it is empty.
@@
This brings us back to the idea behind the proposed definition of a Pareto critical point of a discrete multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$. Let $\sigma\in K$ such that $f(\sigma)=u$. To take the idea of Lipiński et al. [51], the level set $L_{u}$ may be seen as a black box inside of which we know nothing of the dynamics. Instead, to understand the behaviour of $f$ around $\sigma$, we have to look at the boundary of $L_{u}$. Notice that $\operatorname{Ex}L_{u}=\operatorname{Cl}L_{u}\backslash L_{u}\subseteq\operatorname{Cl}L_{u}\backslash\operatorname{Int}L_{u}=\operatorname{Bd}L_{u}$, so the exit set $\operatorname{Ex}L_{u}$ is part of the boundary of $L_{u}$, and from the definition of a multifiltering function, we could show that $\operatorname{Ex}L_{u}=\{\tau\in\operatorname{Cl}L_{u}\ |\ f(\tau)\precneqq f(\sigma)\}$. Hence, $\operatorname{Ex}L_{u}$ corresponds to the part of the domain neighbouring $\sigma$ for which the components of $f$ decrease. Thus, when the relative homology of $\operatorname{Cl}L_{u}$ with respect to $\operatorname{Ex}L_{u}$ is trivial, we interpret that all components of $f$ behave similarly and agree in some common direction around $\sigma$, represented by $\operatorname{Ex}L_{u}$.

### v3

@@ 3. $\mathrm{(3)}$
That said, previous algorithms that generate discrete gradient fields outperform GenerateMDM in terms of speed. In particular, ComputeDiscreteGradient computes a gradient field compatible with the max-extension of a component-wise injective vertex map in $O(|K|+|K_{0}|\log|K_{0}|)$ time, assuming $\dim K$ is low and the star of each vertex is small. However, under the same hypotheses and assuming $k$ is also small, GenerateMDM takes $O(|K|+|f(K)|\log|f(K)|)$ time where $|K_{0}|\leq|f(K)|\leq|K|$. Hence, at best, it can be as efficient as the procedure ComputeDiscreteGradient, but it is slower when the number $|f(K)|$ of level sets of $f$ is great. Nonetheless, as mentioned above, function GenerateMDM proves itself to be advantageous on many other levels.
@@ 5.1. Defining critical simplices for a multifiltering function
Consider a combinatorial gradient field ${\mathcal{V}}$ compatible with a multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$. As stated in Section 3.2, we know that for any $u\in f(K)$, the number $m_{p}(u)$ of critical $p$-simplices of ${\mathcal{V}}$ in $L_{u}$ is at least $\operatorname{rank}H_{p}(K(u),\bigcup_{u^{\prime}\precneqq u}K(u^{\prime}))$. This expression can be rewritten in terms of $L_{u}$. Indeed, notice that $\bigcup_{u^{\prime}\precneqq u}K(u^{\prime})$ is simply $K(u)\backslash L_{u}$. Then, using the excision theorem (see [26]) to excise the set $K(u)\backslash\operatorname{Cl}L_{u}$, we see that
@@
Let ${\mathcal{V}}$ be an acyclic discrete vector field compatible with a multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$. For all $u\in f(K)$, if $H_{*}\big{(}\operatorname{Cl}L_{u},\operatorname{Ex}L_{u}\big{)}$ is nonzero, then ${\mathcal{V}}$ has a critical simplex in $L_{u}$. Conversely, if ${\mathcal{V}}$ is perfect relatively to $f$ and has a critical simplex in $L_{u}$, then $H_{*}\big{(}\operatorname{Cl}L_{u},\operatorname{Ex}L_{u}\big{)}$ is nonzero.
@@
Definition 5.2.
For a multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$, we say that $u\in f(K)$ is a Pareto critical value of $f$ if the relative homology $H_{*}(\operatorname{Cl}L_{u},\operatorname{Ex}L_{u})$ is nonzero. Similarly, $\sigma\in K$ is Pareto critical for $f$ if it belongs in a connected component $C$ of $L_{f(\sigma)}$ such that $H_{*}(\operatorname{Cl}C,\operatorname{Ex}C)$ is nonzero. The Pareto critical simplices of $f$ constitute the Pareto set of $f$, noted $\mathcal{P}_{f}$.
Remark 5.3.
@@
Since the gradient field of a mdm function is a particular case of acyclic discrete vector field, we obtain the following result directly from Proposition 5.1.
Corollary 5.4.
Let $g:K\rightarrow{\mathbb{R}}^{k}$ be a mdm function compatible with a multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$. For every Pareto critical value $u$ of $f$, the mdm function $g$ necessarily has a critical simplex in $L_{u}$. Conversely, when $g$ is perfect relatively to $f$, for every critical simplex $\sigma$ of $g$, we have that $\sigma\in\mathcal{P}_{f}$.
Remark 5.5.
Notice that, when $g$ is not relative-perfect, we may not assume that all critical simplices of $g$ are Pareto for $f$. Indeed, we may find a counterexample by letting $f$ be such that $\mathcal{P}_{f}\neq K$ and $g$ be any mdm function for which all simplices are critical. In that case, $g$ is trivially $f$-compatible, but there exists critical simplices of $g$ which are not Pareto for $f$. That being said, the relative-perfectness hypothesis is not necessary for every critical simplex of $g$ to be in $\mathcal{P}_{f}$. For instance, if $\mathcal{P}_{f}=K$, which is the case notably when $f=0$, then all critical simplices of any $f$-compatible mdm function $g$ trivially belong in $\mathcal{P}_{f}$, whether or not $g$ is relative-perfect.
@@
Our definition is also inspired from the theory of combinatorial multivector fields, which provides discrete analogues to many concepts from smooth dynamics [52]. Indeed, consider a multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$ and some $u\in f(K)$. For any two simplices $\sigma\leq\sigma^{\prime}$, if $f(\sigma)=f(\sigma^{\prime})=u$, then for all $\tau\in K$ such that $\sigma\leq\tau\leq\sigma^{\prime}$, we have $f(\tau)=u$ because $f(\sigma)\preceq f(\tau)\preceq f(\sigma^{\prime})$ by the definition of a multifiltering function. Hence, $\sigma,\sigma^{\prime}\in L_{u}$ implies $\tau\in L_{u}$ whenever $\sigma\leq\tau\leq\sigma^{\prime}$, meaning that $L_{u}$ is convex with respect to the face relation in $K$. In other words, each $L_{u}$ is a combinatorial multivector in the sense of [52]. Therefore, the partition of $K$ in level sets is a well-defined combinatorial multivector field for which each $L_{u}$ is a multivector. Finally, in [52], a multivector $L_{u}$ is defined as critical when $H_{*}(\operatorname{Cl}L_{u},\operatorname{Ex}L_{u})$ is nonzero, so our proposed definition of a Pareto critical value agrees with the combinatorial multivector field theory.
@@
Figure 6. In each figure is shown a neighbourhood $U$ of some point $p$ on the sphere $S^{2}$. For $f:S^{2}\rightarrow{\mathbb{R}}^{2}$ which maps each point to its coordinates $(x,y)\in{\mathbb{R}}^{2}$, the Pareto set of $f$ consists of the two red arcs, while points $q\in U$ such that $f(q)\preceq f(p)$ are orange, those for which $f(p)\preceq f(q)$ are green and all others are yellow. In 6a, the considered point $p$ is regular. In 6b, $p$ is a Pareto minimum and in 6c, it is a Pareto maximum.
@@
To further justify the proposed definition of Pareto singularity, we present concisely the idea of Pareto points for smooth maps and give an interpretation which is common to both discrete and smooth settings. The reader is referred to [47] for more details regarding the basic concepts of differential geometry. Consider a smooth map $f=(f_{1},...,f_{k}):M\rightarrow{\mathbb{R}}^{k}$ on a manifold $M$ and let $p\in M$. We say $p$ is in the Pareto set $\theta$ of $f$ if there exists no direction in which the gradient vectors of all components $f_{i}$ at $p$ agree. In other words, away from $\theta$, all $f_{i}$ increase in some common direction. For all $p\notin\theta$ and for some small enough neighbourhood $U$ of $p$, we could show that this common direction may be represented by the subset $\{q\in U\ |\ f_{i}(q)>f_{i}(p)\text{ for each }i=1,...,k\}$, which is homeomorphic to the intersection of a convex cone with some small neighbourhood, so it is a contractible subset of $U$. An example of such a subset is represented in green in Figure 6a. Generally, this is not true for points $p\in\theta$. For example, in Figure 6b, we see that the subset $\{q\in U\ |\ f_{i}(q)>f_{i}(p)\text{ for each }i=1,...,k\}$ has two distinct components while, in Figure 6c, it is empty.
@@
This brings us back to the idea behind the proposed definition of a Pareto critical point of a discrete multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$. Let $\sigma\in K$ such that $f(\sigma)=u$. As explained in [52], the level set $L_{u}$ may be seen as a black box inside of which we know nothing of the dynamics. Instead, to understand the behaviour of $f$ around $\sigma$, we have to look at the boundary of $L_{u}$. Notice that $\operatorname{Ex}L_{u}=\operatorname{Cl}L_{u}\backslash L_{u}\subseteq\operatorname{Cl}L_{u}\backslash\operatorname{Int}L_{u}=\operatorname{Bd}L_{u}$, so the exit set $\operatorname{Ex}L_{u}$ is part of the boundary of $L_{u}$, and from the definition of a multifiltering function, we could show that $\operatorname{Ex}L_{u}=\{\tau\in\operatorname{Cl}L_{u}\ |\ f(\tau)\precneqq f(\sigma)\}$. Hence, $\operatorname{Ex}L_{u}$ corresponds to the part of the domain neighbouring $\sigma$ for which the components of $f$ decrease. Thus, when the relative homology of $\operatorname{Cl}L_{u}$ with respect to $\operatorname{Ex}L_{u}$ is trivial, we interpret that all components of $f$ behave similarly and agree in some common direction around $\sigma$, represented by $\operatorname{Ex}L_{u}$.